# Collective Noise Filtering in Complex Networks

### v1

@@ Abstract
We focus on the year 2001, which spans the peak of the crisis. After removing self-emails, the dataset contains 12,684 time-stamped emails among 163 employees. We construct the full-year network as a weighted, directed email exchange frequency network, where each directed edge weight is the email frequency from sender to recipient (Figure 4A). The resulting network exhibits a clear core–periphery organization, suggesting complex social structure [20]. The full-year network is a temporal aggregate of a dynamic communication process. To examine temporal variability, we also construct the twelve calendar-month snapshots of 2001 (SI Figure [placeholder]), each of which can be viewed as a noisy sample of the underlying full-year network, where the finite observation window causes sampling fluctuations in the edge weights. An interesting denoising task is therefore to filter each calendar-month snapshot to bring it closer to the full-year network, which we use as a reference signal.
@@
To address the first goal, we use month-to-month fluctuations to estimate edge-level noise covariance and use this information in the NetWF, producing twelve denoised calendar-month snapshots (SI Figure [placeholder]). For the OS baseline, only the mean variance over nodes extracted from the full noise covariance is used (SI Figure [placeholder]). October 2001 provides an illustrative example because of its historical significance: Enron’s accounting fraud came to light, sharply accelerating the company’s downfall. Relative to the full-year network, the raw October snapshot shows turmoil with pronounced sampling fluctuations, including the presence of unusually strong edges (Figure 4B). After applying the NetWF, these extremes are smoothed while the global structure remains preserved (Figure 4C). Globally, in terms of the MSE with respect to the full-year network, the NetWF again outperforms both the raw snapshot and the OS. Over all twelve months, the NetWF achieves an average MSE of $0.11\pm 0.01$, lower and more robust compared to the raw snapshots ($0.21\pm 0.04$) and the OS outcomes ($0.17\pm 0.03)$, as seen in Figure 4D. Similar evaluations carried out for each individual calendar-month snapshot, including both MSE and $R^{2}$ perspectives, are shown in SI Figure [placeholder]. There, we also test the restricted version of the NetWF that uses homogenized noise information, in line with the OS. Paired $t$-tests show that the full NetWF reduces MSE significantly relative to both the raw snapshots ($p=0.01$) and the OS ($p=0.03$), while the second-best restricted NetWF is still significantly better than the raw snapshots ($p=0.02$).
@@
Upon further testing the NetWF, we also found an interesting directional asymmetry in this network. If we modify Eq. 5 by using source-node similarity in place of target-node similarity, i.e. defining $S_{\,\overrightarrow{AB},\,\overrightarrow{CD}}\equiv s^{\text{source}}_{A,C}\,s^{\text{source}}_{B,D}$, the NetWF delivers even stronger denoising results with an average MSE of $0.07\pm 0.01$, as shown in SI Figure [placeholder]. This suggests that similarity in sending behaviors might carry more information about the edge weight correlations. This example illustrates that although our proposed edge-similarity measure is generically applicable, even better performance can be achieved if the similarity measure is tailored to the structure of the specific network dataset.
@@
Last, we turn to our second goal with this dataset, that is, denoising when $C_{n}$ is unknown. In this setting, we use a naïve proxy of noise level for both the NetWF and the OS, using the variance across edge weights in the observed single calendar-month snapshot as the homogeneous noise variance. Even under this under-specified condition, the NetWF achieves significantly lower average MSE ($0.14\pm 0.01$) compared to both the raw snapshots and the OS baseline ($0.18\pm 0.03$) according to paired $t$-tests, with $p$-values $0.02$ and $0.04$ respectively. Therefore, successful denoising is possible with the NetWF even with minimal inputs, as long as a reasonable noise level is assumed; see SI Figure [placeholder].
@@ Discussion
The Wiener filter was the first example of a statistically solid noise reduction technique [61]. The generalized Wiener filtering framework is mathematically grounded, as the Wiener operator (Eq. 2) optimally solves the MSE minimization problem (Eq. 1). Yet, an extension to even static network data remained an unsolved problem until now. Our proposed NetWF offers a qualitative advancement in addressing the often-overlooked issue of noise in static networks. It is a versatile denoising scheme that is applicable to both directed and undirected networks, with any weights, including signed networks. A key to this advancement is our proposed concept of edge-similarity, a generic solution that is intuitive and interpretable, building upon any preferred notion of node similarity.
@@
An exciting future direction is to utilize the NetWF to identify the most suitable measure of node (or edge) similarity adaptively in each dataset that optimizes the output of the NetWF. As illustrated on the Enron email dataset, such notions of similarity that are tailored to each application are expected to lead to further improved results. Another promising avenue is to use dimension reduction and machine learning techniques to learn the optimal similarity structure, such as using node2vec to generate low-dimensional node embeddings, which can then be used to model node and edge-similarities [27]. The NetWF naturally adapts to different levels of uncertainty in network data, see SI Figure [placeholder]. Although the NetWF is expected to perform best when a full noise characterization is available, we illustrated that in practice it is able to yield similarly good results even with much less information, such as a homogeneous noise estimate.
@@
As large-scale biological data are inherently imperfect, this is a key area of application for the NetWF. As an illustration, we used a biologically coherent portion of the entire SGA GI dataset in the baker’s yeast, where the strong reported structure of the data offered multiple reliable evaluations by comparing to benchmark datasets. A natural next step is to apply the NetWF to the entire yeast data as well as to similar data in other organisms, like the recent genome-scale screen in human HAP1 cells [12]. By reducing data noise, these applications could further clarify the amount of agreement between the functional architecture of various organisms.
@@
As illustrated on the Enron dataset, the static NetWF is also applicable to snapshots over temporal networks. An opportunity for future development is a temporal NetWF that considers the full time series of network data, combining temporal Wiener filtering ideas with the NetWF. Further extensions include multiplex and multilayer networks [13, 30], as well as hypergraphs and simplicial complexes [8, 11]. Such higher-order interactions play an important role in genetics, as evidenced by the prevalence of GIs between sets of three genes in yeast [35, 34], and are also increasingly studied in computational social science [53, 40].
@@
Just as with any form of quantitative data analysis, understanding and reporting uncertainties in networks is crucial, as these uncertainties can be leveraged to improve the overall quality and reliability of network-based inference. We believe that the NetWF is an important starting point towards a paradigm change in network science where reporting and utilizing noise characteristics for any network data is a valuable and integral part of understanding the underlying complex networks.
@@
We consider an observed network on $v$ nodes, represented by a demeaned adjacency matrix $A\in\mathbb{R}^{v\times v}$ and its vectorization $a=\mathrm{vec}(A)\in\mathbb{R}^{k}$ with $k=v^{2}$, together with a noise covariance matrix $C_{n}\in\mathbb{R}^{k\times k}$ characterizing noise in the edge weights. We first compute node-level profile similarity networks (PSNs), symmetric matrices whose elements are similarities between node pairs, as in Eqs. 34 for directed networks or Eq. 6 for undirected networks. We present two NetWF implementations: a direct NetWF and an iterative NetWF.
@@
Hence, we develop an alternative, iterative implementation of the NetWF, which reduces the memory burden to $O(v^{2})$, as long as $C_{n}$ is diagonal (i.e. noise is treated as uncorrelated across edges). The key is to avoid explicitly forming $C_{u}$. Since $C_{u}+C_{n}$ is symmetric and positive semidefinite, we deploy the conjugate gradient (CG) algorithm to efficiently solve the linear subproblem $(C_{u}+C_{n})x=a$ with guaranteed numerical convergence [26]. Note that CG does not require explicit formation of $C_{u}+C_{n}$, but only its action on vectors. Consequently, in both the CG subproblem and the subsequent $C_{u}x$ computation, we exploit the Kronecker-product structure of $C_{u}$ via the identity $(M_{1}\otimes M_{2})x=\mathrm{vec}\left(M_{2}\,\mathrm{unvec}(x)\,M_{1}^{\mathsf{T}}\right)$, reducing the dimensionality of the problem from $\mathbb{R}^{k\times k}$ to $\mathbb{R}^{v\times v}$. This step resolves the memory bottleneck and allows the NetWF to scale to large empirical systems.
Pseudocode for both the direct NetWF and the CG-based iterative NetWF are provided in SI Text [placeholder].
@@
The S. cerevisiae GI network, along with the noise variance matrix, was constructed from the SGA results reported in Ref. [19]. We focus on the densely connected essential GI network (ExE), in which essential genes are perturbed using temperature-sensitive alleles. Because each gene pair may be assayed in two orientations (query vs. array), we symmetrize the data by averaging both the reported GI scores and their associated noise variances, resulting in an undirected, weighted network on $427$ essential-gene nodes.
Note that the raw ExE matrix, as well as the corresponding variances, contains missing entries; in cross validation, we likewise mask entries to create additional missing values. In the cross validation, we also manually put in missing entries. Before ingesting the data into the NetWF or the OS, we impute missing noise variances with the sum of the mean of observed variances and the variance of observed GI scores, and then impute missing GI scores with the mean of observed GI scores.
The PPI benchmark network was obtained from BioGRID [48], version 5.0.253 (compiled on Dec. 25, 2025), at https://thebiogrid.org. We filtered interaction types to retain only physical association and direct interaction. The protein co-complex membership benchmark network was derived from the protein complex standard provided by Ref. [19] based on Refs. [6, 10]. The Gene Ontology (GO) co-annotation network was constructed following the protocol in Ref. [19], with the GO functional annotation data downloaded from the Saccharomyces Genome Database (SGD) [17] at https://www.yeastgenome.org.
The Enron email corpus was obtained from https://www.cs.cmu.edu/~enron/, using the latest May 7, 2015 release. Both the NetWF and the OS baseline can introduce self-links and negative edge weights, even when the input network contains no self-links or negative edge weights. For visualization and evaluation in the Enron application, we therefore remove self-links and truncate negative weights to zero for all denoised outputs.
@@
For full reproducibility of the NetWF results in this study, we provide relevant processed data and code at https://github.com/markzhao98/NetWF_paper. – – – – – – – – – –
@@ Acknowledgments
The authors acknowledge funding through the ‘CAREER: Network-based inference of complex biological interactions’ PHY-2440223 Physics of Living Systems (POLS) NSF CAREER Award, sponsored by the NSF 22-586 Faculty Early Career Development Program. We thank Bingjie Hao, Ruiting Xie, Anastasiya Salova, Eli Daniel Ganz, Maryn Carlson, Noshir Contractor, Michael Costanzo, and Charles Boone for their helpful comments and discussion. T.Z. and I.A.K. acknowledge support from the NSF–Simons National Institute for Theory and Mathematics in Biology, jointly funded by the U.S. National Science Foundation (Award No. 2235451) and the Simons Foundation (Award No. MP-TMPS-00005320).

### v2

@@ Abstract
We focus on the year 2001, which spans the peak of the crisis. After removing self-emails, the dataset contains 12,684 time-stamped emails among 163 employees. We construct the full-year network as a weighted, directed email exchange frequency network, where each directed edge weight is the email frequency from sender to recipient (Figure 4A). The resulting network exhibits a clear core–periphery organization, suggesting complex social structure [20]. The full-year network is a temporal aggregate of a dynamic communication process. To examine temporal variability, we also construct the twelve calendar-month snapshots of 2001 (SI Figure 5), each of which can be viewed as a noisy sample of the underlying full-year network, where the finite observation window causes sampling fluctuations in edge weights. An interesting denoising task is therefore to filter each calendar-month snapshot to see if that brings it closer to the full-year network, which we use as a reference signal.
@@
To address the first goal, we use month-to-month fluctuations to estimate edge-level noise covariance and incorporate this information into the NetWF, producing twelve denoised calendar-month snapshots (SI Figure 6). For the OS baseline, only the mean variance over nodes calculated from the full noise covariance is used (SI Figure 7). October 2001 provides an illustrative example because of its historical significance: Enron’s accounting fraud came to light, sharply accelerating the company’s downfall. Relative to the full-year network, the raw October snapshot shows turmoil with pronounced sampling fluctuations, including the presence of unusually strong edges (Figure 4B). After applying the NetWF, these extremes are smoothed while the overall structure remains preserved (Figure 4C). Globally, in terms of the MSE with respect to the full-year network, the NetWF again outperforms both the raw snapshot and the OS. Over all twelve months, the NetWF achieves an average MSE of $0.11\pm 0.01$, lower and more robust compared to the raw snapshots ($0.21\pm 0.04$) and the OS outcomes ($0.17\pm 0.03)$, as seen in Figure 4D. Similar evaluations carried out for each individual calendar-month snapshot, including both MSE and $R^{2}$ perspectives, are shown in SI Figure 8. There, we also test the restricted version of the NetWF that uses homogenized noise information, in line with the OS. Paired $t$-tests show that the full NetWF reduces MSE significantly relative to both the raw snapshots ($p=0.01$) and the OS ($p=0.03$). The restricted NetWF is still significantly better than the raw snapshots ($p=0.02$), although its improvement over the OS is not significant ($p=0.10$).
@@
Upon further testing the NetWF, we also found an interesting directional asymmetry in this network. If we modify Eq. 5 by using source-node similarity in place of target-node similarity, i.e. defining $S_{\,\overrightarrow{AB},\,\overrightarrow{CD}}\equiv s^{\text{source}}_{A,C}\,s^{\text{source}}_{B,D}$, the NetWF delivers even stronger denoising results with an average MSE of $0.07\pm 0.01$, as shown in SI Figure 9. This suggests that similarity in sending behaviors might carry more information about the edge weight correlations in this network. More broadly, this finding highlights that although our proposed edge-similarity measure is generically applicable, even better performance can be achieved if the similarity measure is tailored to the structure of the specific network dataset.
@@
Last, we turn to our second goal with this dataset, that is, denoising when $C_{n}$ is unknown. In this setting, we use a naïve guess of noise level for both the NetWF and the OS, using the variance across edge weights in the observed single calendar-month snapshot as the homogeneous noise variance. Even under this under-informed condition, the NetWF achieves significantly lower average MSE ($0.14\pm 0.01$) compared to both the raw snapshots and the OS baseline ($0.18\pm 0.03$) according to paired $t$-tests, with $p$-values $0.02$ and $0.04$ respectively. Therefore, successful denoising is possible with the NetWF even with minimal inputs, as long as a reasonable noise level is assumed; see SI Figure 10.
@@ Discussion
Wiener filtering was the first example of a statistically solid, collective noise reduction technique [61]. The generalized Wiener filter framework is mathematically grounded, as the Wiener operator (Eq. 2) optimally solves the MSE minimization problem (Eq. 1). Yet, an extension to even static network data remained an unsolved problem until now. Our proposed NetWF offers a qualitative advancement in addressing the often-overlooked issue of noise in static networks. It is a versatile denoising method applicable to both directed and undirected networks, with arbitrary weights, including signed networks. A key to this advancement is our proposed concept of edge similarity, a generic solution that is intuitive and interpretable, building upon any preferred notion of node similarity.
@@
An exciting future direction is to utilize the NetWF to identify the most suitable measure of node (or edge) similarity adaptively in each dataset that optimizes the output of the NetWF. As investigated in our Enron study, such notions of similarity that are tailored to each application domain are expected to lead to further improved results. Another promising avenue is to use dimension reduction and machine learning techniques to learn the optimal similarity structure, such as using node2vec to generate low-dimensional node embeddings, which can then be used to model node (and edge) similarities [27]. The NetWF naturally adapts to different levels of uncertainty in network data; see SI Figure 11. Although the NetWF is expected to perform best when a full noise characterization is available, we have seen that it can deliver similarly good results even with much less information, such as a homogeneous noise estimate.
@@
As large-scale biological data are inherently imperfect, this is a key area of application for the NetWF. As an illustration, we used a biologically coherent ExE portion of the entire SGA GI dataset for baker’s yeast, where the strong structure of the network offered multiple reliable evaluations against benchmark datasets. A natural next step is to apply the NetWF to the entire yeast dataset as well as to similar datasets in other organisms, such as the recent genome-scale screen in human HAP1 cells [12]. By reducing data noise, these applications could further clarify the amount of agreement in the functional architecture across various organisms.
@@
As showcased with the Enron Corpus email network study, the static NetWF is also applicable to snapshots of temporal networks. An opportunity for future development is a temporal NetWF that considers the full time series of network data, combining temporal Wiener filtering ideas with the NetWF. Further extensions include multiplex and multilayer networks [13, 30], as well as hypergraphs and simplicial complexes [8, 11]. Such higher-order interactions play an important role in genetics, as evidenced by the prevalence of GIs between sets of three genes in yeast [35, 34], and are also increasingly studied in computational social science [53, 40].
@@
Just like in any form of quantitative data analysis, understanding and reporting uncertainties in networks is crucial, as these uncertainties can be leveraged to improve the overall quality and reliability of network-based inference. We believe that the NetWF serves as an important milestone towards a paradigm shift in network science, where reporting and utilizing noise characteristics for network data is seen as a valuable and integral part of understanding the underlying complex networks.
@@
We consider an observed network on $v$ nodes, represented by a demeaned adjacency matrix $A\in\mathbb{R}^{v\times v}$ and its vectorization $a=\mathrm{vec}(A)\in\mathbb{R}^{k}$ with $k=v^{2}$, together with a noise covariance matrix $C_{n}\in\mathbb{R}^{k\times k}$ characterizing noise in edge weights. We first compute node-level profile similarity networks (PSNs), symmetric matrices whose elements are similarities between node pairs, as in Eqs. 34 for directed networks or Eq. 6 for undirected networks. We present two NetWF implementations: a direct NetWF and an iterative NetWF.
@@
Hence, we develop an alternative, iterative implementation of the NetWF, which reduces the memory burden to $O(v^{2})$, as long as $C_{n}$ is diagonal (i.e. noise is treated as uncorrelated across edges). The key is to avoid explicitly forming $C_{u}$. Since $C_{u}+C_{n}$ is symmetric and positive semidefinite, we deploy the conjugate gradient (CG) algorithm to efficiently solve the linear subproblem $(C_{u}+C_{n}+\epsilon I)x=a$ with guaranteed numerical convergence [26]. Note that CG does not require explicit formation of $C_{u}+C_{n}$, but only its action on vectors. Consequently, in both the CG subproblem and the subsequent $C_{u}x$ computation, we exploit the Kronecker-product structure of $C_{u}$ via the identity $(M_{1}\otimes M_{2})x=\mathrm{vec}\left(M_{2}\,\mathrm{unvec}(x)\,M_{1}^{\mathsf{T}}\right)$, reducing the dimensionality of the problem from $\mathbb{R}^{k\times k}$ to $\mathbb{R}^{v\times v}$. This step resolves the memory bottleneck and allows the NetWF to scale to large empirical systems.
@@
The S. cerevisiae GI network and its associated noise variance matrix were constructed from the SGA dataset reported in Ref. [19] and made available at https://thecellmap.org/yeast/costanzo2016/. We study the densely connected essential GI network (ExE), between essential genes. The dataset provides GI scores, double-mutant fitness standard deviations, and $p$-values for assays between 1,1081{,}108 query strains (covering $796$ query genes) and $792$ array strains (covering $561$ array genes). For query–array gene pairs with duplicated assays, we select the assay with the most significant $p$-value and drop the others. We then take the union of unique query and array genes, yielding $855$ essential genes, and assemble a square GI matrix indexed by this gene set. If two experiments map to the same (undirected) gene pair, we average their $\epsilon$ values. This procedure produces a symmetric GI matrix representing an undirected, weighted network on the $855$ essential gene nodes. The noise variance matrix is constructed analogously: for each assay we square the reported double-mutant fitness standard deviation $\sigma$ to obtain an edge-level variance $\sigma^{2}$, place this value into the same $855\times 855$ gene-indexed matrix, and, when multiple assays map to the same (undirected) gene pair, average their $\sigma^{2}$ values.
Note that the raw ExE matrix, as well as the corresponding variances, contains missing entries. During cross validation, we mask more entries to create additional missing values. Before ingesting the data into the NetWF (including PSN computation) or the OS, we impute missing noise variances with the sum of the mean of observed variances over entries in the noise variance matrix and the variance of observed GI scores over entries in the ExE matrix, and then impute missing GI scores with the mean of observed GI scores over entries in the ExE matrix.
The PPI benchmark network was obtained from BioGRID [48], version 5.0.253 (compiled on Dec. 25, 2025), at https://thebiogrid.org, filtering interaction types to retain only physical association and direct interaction. The protein co-complex membership benchmark network was derived from the protein complex standard provided by Ref. [19] based on Refs. [6, 10]. The Gene Ontology (GO) bioprocess term co-annotation benchmark network was constructed following the protocol in Ref. [19], with the GO bioprocess term functional annotation data downloaded from the Saccharomyces Genome Database (SGD) [17] at https://www.yeastgenome.org.
The Enron Corpus email network was obtained from https://www.cs.cmu.edu/~enron/, using the latest May 7, 2015 release. Both the NetWF and the OS baseline can introduce self-links and negative edge weights, even when the input network contains no self-links or negative edge weights. For visualization and evaluation in the Enron application, we therefore remove self-links and truncate negative weights to zero for all denoised outputs.
@@
For full reproducibility of the NetWF results in this study, an archived release of relevant code and processed data is available on Zenodo: 10.5281/zenodo.18472712. The development repository is maintained at https://github.com/markzhao98/NetWF_paper.
@@ Acknowledgments
The authors acknowledge funding through the ‘CAREER: Network-based inference of complex biological interactions’ PHY-2440223 Physics of Living Systems (POLS) NSF CAREER Award, sponsored by the NSF 22-586 Faculty Early Career Development Program. We thank Bingjie Hao, Ruiting Xie, Anastasiya Salova, Eli Daniel Ganz, Leone Luzzatto, Maryn Carlson, Noshir Contractor, Michael Costanzo, and Charles Boone for their helpful comments and discussion. T.Z. and I.A.K. acknowledge support from the NSF–Simons National Institute for Theory and Mathematics in Biology, jointly funded by the U.S. National Science Foundation (Award No. 2235451) and the Simons Foundation (Award No. MP-TMPS-00005320).